# The formation and evolution of dust in the colliding-wind binary Apep revealed by JWST

### v1

@@ 6.4 Distance to Apep
While an anisotropic wind velocity field has been suggested to explain the difference between the sky-plane dust expansion speed and line-of-sight wind speed (see Section 6.1.3, J. R. Callingham et al., 2019), an alternative possibility is that the distance to Apep is significantly larger than previously estimated. If Apep’s distance is underestimated by a factor of 2 to 3, this would bring the system in line with other WR binaries for which the two measurements are approximately consistent (N. Richardson, 2025). The distance of Apep has been difficult to determine from Gaia astrometry due to its binary nature (J. R. Callingham et al., 2019). Apep has also not been identified as a member of an OB association (J. R. Callingham et al., 2019). Dropping the 2.4 kpc assumption used so far in this study, the temperature profile, which constrains the stellar luminosity as a function of radius in au terms, offers constraints on the distance from a thermal perspective.
@@
Importantly, this is different from the dependence between $L$ and $d$ provided by flux density constraints, in which case $L$ scales with $d^{2}$ to maintain a given flux density for the system. J. R. Callingham et al. (2019) used the V-band apparent magnitude to place constrains on the distance based on typical WR absolute magnitudes. Here, we use the near-IR SED (Y. Han et al., 2020) of the WR binary and the northern companion in the system to calibrate the scaling between luminosity and distance. Using the same WC:WN:O luminosity ratio and Potsdam WR and O star spectra as assumed in Section 6.2 to maintain their near-IR flux density ratio, the required scaling between the luminosity of the WR binary and the O star companion is shown in Fig. 7.
@@
The different dependence between luminosity and distance required under the thermal profile and photometric constraints demands that the two constraints cross at only one point (apart from the origin) within the luminosity and distance parameter space, which in this case is 4.6 kpc. This is almost twice the value from previous estimates (J. R. Callingham et al., 2019; Y. Han et al., 2020). In practice, uncertainties in both constraints relax the point of intersection into a region of parameter space over a wide interval of plausible distances. However, further distance constraints are provided by the dynamics the binary’s orbit. Given the orbital parameters derived for the system based on the dust geometry (R. White et al., submitted), the present position angle of the WR binary determined with near-IR interferometry (Y. Han et al., 2020) and radio imaging of the colliding-wind region (B. Marcote et al., 2021) places it near apastron, translating the $47\pm 6$ mas projected separation (Y. Han et al., 2020) to a semi-major axis of $28_{-6}^{+10}$ mas. Given the 193 yr orbital period that we find, only a limited range of distances can permit such a semi-major axis for typical WC+WN masses between 20 and 40 M (P. A. Crowther, 2007; A. A. C. Sander et al., 2019; Y. Han et al., 2020), which is indicated in Fig. 7. The intersection between the thermal, photometric and dynamical constraints limit the range of plausible distances to between 2.8 to 5.1 kpc, which also sits comfortably within the typical range of combined WN and WC luminosities given the spectral classification of Apep (W. R. Hamann et al., 2006), with the northern companion also at luminosities comparable to other stars of its spectral type (F. Martins et al., 2005; P. A. Crowther & C. J. Evans, 2009). Note that there could be additional lines of evidence related to Apep’s distance that are not considered in depth here. These include diffuse interstellar bands (DIBs) present in the spectra of the system, which suggest an upper limit of 4.5 kpc (G. Zasowski & B. Ménard, 2014; J. R. Callingham et al., 2019); the radio luminosity of Apep, which would be similar to the baseline level (i.e., not during outburst) of the well-studied colliding-wind binary $\eta$ Car (K. Davidson & R. M. Humphreys, 1997) at this upper limit from DIBs (J. R. Callingham et al., 2019); and the X-ray luminosity of Apep, which would be among the most luminous colling-wind binary known at this distance (J. R. Callingham et al., 2019).
@@ 7 Conclusions
    We determined the temperature of dust as a function of radius by modelling the spatially resolved SED of the dust structure assembled from the multi-wavelength observations. The temperature profile that we measured is largely consistent with the optical properties of amorphous carbon dust grains.
@@
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #5842. Support for program #5842 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. RMTW acknowledges the financial support of the Andy Thomas Space Foundation. BJSP was funded by the Australian Government through the Australian Research Council DECRA fellowship DE210101639. NDR is grateful for support from the Cottrell Scholar Award #CS-CSA-2023-143 sponsored by the Research Corporation for Science Advancement. {contribution} YH planned the JWST and VLT observations, carried out data reduction and analyses and wrote the manuscript. All authors contributed to discussing scientific interpretations and editing the manuscript.
@@
The JWST/MIRI data used in this study is available on the Mikulski Archive for Space Telescopes under JWST program ID 5842. The VLT/VISIR data used in this study are available on the ESO archive under program IDs 097.C-0679, 0101.C-0726 and 113.26A7.001.
@@ Appendix B Resolved SED fitting
Figure 10: The spatially resolved SED of Apep constructed from a range of shell orders and observing dates. The SED models of each feature at each epoch is shown with dotted lines. The top, middle and bottom rows correspond to the E1, E2 and E3 features respectively.

### v2

@@ 6.4 Distance to Apep
While an anisotropic wind velocity field has been suggested to explain the difference between the sky-plane dust expansion speed and line-of-sight wind speed (see Section 6.1.3, Callingham2019), an alternative possibility is that the distance to Apep is significantly larger than previously estimated. If Apep’s distance is underestimated by a factor of 2 to 3, this would bring the system in line with other WR binaries for which the two measurements are approximately consistent (Richardson2025). The distance of Apep has been difficult to determine from Gaia astrometry due to its binary nature (Callingham2019). Apep has also not been identified as a member of an OB association (Callingham2019). Dropping the 2.4 kpc assumption used so far in this study, the temperature profile, which constrains the stellar luminosity as a function of radius in au terms, offers constraints on the distance from a thermal perspective.
@@
Importantly, this is different from the dependence between $L$ and $d$ provided by flux density constraints, in which case $L$ scales with $d^{2}$ to maintain a given flux density for the system. Callingham2019 used the V-band apparent magnitude to place constrains on the distance based on typical WR absolute magnitudes. Here, we use the near-IR SED (Han2020) of the WR binary and the northern companion in the system to calibrate the scaling between luminosity and distance. Using the same WC:WN:O luminosity ratio and Potsdam WR and O star spectra as assumed in Section 6.2 to maintain their near-IR flux density ratio, the required scaling between the luminosity of the WR binary and the O star companion is shown in Fig. 7.
@@
The different dependence between luminosity and distance required under the thermal profile and photometric constraints demands that the two constraints cross at only one point (apart from the origin) within the luminosity and distance parameter space, which in this case is 4.6 kpc. This is almost twice the value from previous estimates (Callingham2019; Han2020). In practice, uncertainties in both constraints relax the point of intersection into a region of parameter space over a wide interval of plausible distances. However, further distance constraints are provided by the dynamics the binary’s orbit. Given the orbital parameters derived for the system based on the dust geometry (White2025), the present position angle of the WR binary determined with near-IR interferometry (Han2020) and radio imaging of the colliding-wind region (Marcote2021) places it near apastron, translating the $47\pm 6$ mas projected separation (Han2020) to a semi-major axis of $28_{-6}^{+10}$ mas. Given the 193 yr orbital period that we find, only a limited range of distances can permit such a semi-major axis for typical WC+WN masses between 20 and 40 M (Crowther2007; Sander2019; Han2020), which is indicated in Fig. 7. The intersection between the thermal, photometric and dynamical constraints limit the range of plausible distances to between 2.8 to 5.1 kpc, which also sits comfortably within the typical range of combined WN and WC luminosities given the spectral classification of Apep (Hamann2006), with the northern companion also at luminosities comparable to other stars of its spectral type (Martins2005; Crowther2009). Note that there could be additional lines of evidence related to Apep’s distance that are not considered in depth here. These include diffuse interstellar bands (DIBs) present in the spectra of the system, which suggest an upper limit of 4.5 kpc (Zasowski2014; Callingham2019); the radio luminosity of Apep, which would be similar to the baseline level (i.e., not during outburst) of the well-studied colliding-wind binary $\eta$ Car (Davidson1997) at this upper limit from DIBs (Callingham2019); and the X-ray luminosity of Apep, which would be among the most luminous colling-wind binary known at this distance (Callingham2019).
@@ 7 Conclusions
    We determined the temperature of dust as a function of radius by modelling the spatially resolved SED of the dust structure assembled from the multi-wavelength observations. Assuming amorphous carbon grains, the temperature profile that we measured is largely consistent with thermal equilibrium.
@@
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #5842. Support for program #5842 was provided by NASA through a grant from the Space Telescope Science Institute. YH is funded by a Caltech Barr Fellowship. RMTW acknowledges the financial support of the Andy Thomas Space Foundation. BJSP was funded by the Australian Government through the Australian Research Council DECRA fellowship DE210101639. NDR is grateful for support from the Cottrell Scholar Award #CS-CSA-2023-143 sponsored by the Research Corporation for Science Advancement. {contribution}YH planned the JWST and VLT observations, carried out data reduction and analyses and wrote the manuscript. All authors contributed to discussing scientific interpretations and editing the manuscript.
@@
The JWST/MIRI data used in this study is available on the Mikulski Archive for Space Telescopes under JWST program ID 5842 or at https://dx.doi.org/10.17909/mf14-ga19 (catalog doi:10.17909/mf14-ga19). The VLT/VISIR data used in this study are available on the ESO archive under program IDs 097.C-0679, 0101.C-0726 and 113.26A7.001.
@@ Appendix B Resolved SED fitting
Figure 10: The spatially resolved SED of Apep constructed from a range of shell orders and observing dates. The top, middle and bottom rows correspond to the E1, E2 and E3 features respectively. The SED models fitted individually to each feature at each epoch (Section 5.2) are shown with dotted lines, whereas the SED models assuming the best-fit power-law temperature profile (Eq. 4) are shown with dashed lines.